2. Convert the following units. Make sure that your answer has the proper number of significant figures a. 4.8 g/cma to kg/m b. 941 μm to km. Use scientific notation for your answer. (μ is the prefix micro and 1 μm 1 x 10-6 m) c. The speed of light is 2.9989 x 108 m/s. Convert to miles/hour. Use scientific notation in your answer. d. 782 kg to ug. Use scientific notion for your answer.show work please
Answer #1

2.

(a) 941 g/cm3 to kg/m3

We know that , 1kg = 1000 g , so 1 g = (1/1000) kg

And 1 m = 100 cm , so cm = (1/100) m

Now,

1 g/cm3 =[ (1/1000)÷(1/100)3] kg/m3 = 1000 kg/m3

Hence 4.8 g/cm3 = 4.8×1000 kg/m3 = 4800 kg/m3

Answer : 4800 kg/m3

(b) 941 \mum to km

We know that , 1 \mum = 1×10-6 m

Or we can say , 1 m = 106\mum

And 1 km = 1×103 m

Or we can say , 1 m = 10-3 km

So , 1 \mum =( 1×10-6×10-3)km = 1×10-9 km

Hence , 941 \mum = 941×10-9 km = 9.41×10-7 km

Answer : 9.41 ×10-7 km

(c) 2.9989×108 m/s to miles/hour

We know that , 1 m = 6.2137 ×10-4 miles

And 1 s = (1/3600) hour

So 1 m/s = (6.2137×10-4)÷(1/3600) miles/hour

1 m/s = 2.2369 miles/hour

Hence 2.9989×108 m/s = (2.9989×108×2.2369) miles per hour

= 6.7082 × 108 miles/hour

Answer : 6.7082 ×108 miles/hour

(d) 782 kg to \mug

We know that, 1 \mug = 1×10-6 g

And 1 kg = 1×103 g

Or we can say , 1 g = 1×10-3 kg

So 1 kg = 1×109\mug

Hence 782 kg = 782×109\mug = 7.82×1011\mug

Answer: 7.82×1011\mug